Determine the equation of the line
Passing through (5,8)and perpendicular to the line with the equation 5x-2y=10

Answer by ewatrrr(24785)

 
Hi,


Passing through (5,8)and perpendicular to the line with the equation 5x-2y=10
perpendicular to the line with the equation
5x-2y=10 OR y = (5/2)x - 5 m = (5/2)
Perpendicular lines have slopes that are negative reciprocals on one another
New Line
y = (-2/5)x + b |Using ordered pair (5,8) to solve for b
8 = (-2/5)*5 + b
10 = b
y = (-2/5)x + 10 OR 2x + 5y = 50